# Design of a CsI(Tl) Calorimeter for Muonium-to-Antimuonium Conversion Experiment

### v2

@@ II.3.2 Beam Background
A parallel simulation involving $10^{11}$ $e^{+}$ events was conducted for comparison with the signal. No coincidence events were observed in this simulation. According to the technical specifications of the Experimental Muon Source (EMuS) proposed at CSNS [50], the optimal ratio of positrons to muons is expected to be 0.1% [51]. The beam positron background level is estimated based on this data. Assuming a $\mu^{+}$ beam flux of $10^{8}/$s, and a coincidence probability of $10^{-4}$ with the magnetic spectrometer, the upper limit of the background rates is established at 0.007/0.007/0.007 /year at 90% C.L. Thus, the ECAL is expected to be tolerant of both continuous and pulsed beams.
II.3.3 Cosmic-Ray Muon Background
Cosmic-ray muons are another source of background in ECAL. The EcoMug generator [52] is integrated into the MACE offline software to simulate cosmic-ray muon events on various types of surfaces (plane, cylinder, and half-sphere) with accurate angle and momentum distributions.
To evaluate cosmic-ray muon background level in a potential future environment for the MACE detector, a hypothetical 18 m deep tunnel has been modeled in Geant4. A concrete volume, represented by G4Box with dimensions $50\times 50\times 18~{}\text{m}^{3}$, is created, with an air-filled cylindrical region, represented by G4Tubs, subtracted from its bottom to form a tunnel with a 1 m radius and 10 m length. ECAL is positioned at the center of the tunnel, surrounded by a copper solenoid and an iron shield. Cosmic-ray muons are emitted from a $50\times 50~{}\text{m}^{2}$ plane source at sea level. With a rate of 129 Hz, cosmic-ray muon events equivalent to 1 hour of exposure are generated.
The simulation results indicate that cosmic-ray muons still contribute to the background in the ECAL even after passing through the concrete and shielding. This includes some events where $e^{+}e^{-}$ annihilation produces 511 keV $\gamma$ rays. After applying coincidence selection, no events remain, resulting in background rates of less than 2/2/2 /year when considering coincidences with the Magnetic Spectrometer. Typically, a veto system can be installed outside ECAL to further reject cosmic-ray muon background [6, 53].
@@ III.1 Experiment Setups and Results
To characterize the parameters of the CsI(Tl) scintillator, a benchmark detector was constructed and calibrated using various radioactive sources at Sun Yat-sen University. It consists of a $3\times 3\times 8~{}\text{cm}^{3}$ CsI(Tl) crystal and a 1.5” PMT (HAMAMATSU CR284 with E2183-500 socket assembly). The crystal is wrapped with ESR (glued) and coupled with PMT using optical grease, finally coated with 3M tape (Fig. 6b). The readout system uses a Lecroy HDO4054A oscilloscope (Fig. 6c) for signal triggering, integration, and digitization.
@@
(c) Lecroy HDO4054A with a typical waveform of CsI(Tl) on screen.
@@
Figure 7: Measured spectra of ${}^{137}\text{Cs}$ (in green), ${}^{152}\text{Eu}$ (in red) at SYSU and ${}^{22}\text{Na}$ (in blue) at IMP.
The benchmark detector is calibrated using ${}^{137}\text{Cs}$ and ${}^{152}\text{Eu}$ radioactive sources. Radioactive sources were positioned close to the detector, and data acquisition was performed for 1800 s. When the PMT pulse signal exceeds the threshold, the oscilloscope triggers, integrates the waveform and records the charge value. The calibration used the 662 keV peak from ${}^{137}\text{Cs}$ and the 40 keV, 122 keV, 244 keV, 344 keV, and 1408 keV peaks from ${}^{152}\text{Eu}$, as shown in Fig. 7. The spectra display the full-energy peaks of both isotopes. Using the FWHM curve, the energy resolution at 511 keV is calculated to be 11.5%.
A more precise experiment was conducted to investigate the energy resolution at the Institute of Modern Physics, Chinese Academy of Sciences in Lanzhou, China. The absolute gain of the CR284 PMT was accurately calibrated in advance using single photoelectron (SPE) response measurements. A waveform generator sends pulses with a voltage of 2.65 V, a frequency of 10 kHz, and a width of 5 ns to an LED to produce weak light emission. A 120 s measurement scans the high voltage ranging from 850 V to 1250 V. The PMT response waveforms are read out and recorded in ROOT files using the CAEN V1730SB digitizer (Fig. 6d), which is externally triggered by the waveform generator. After offline pulse processing, the SPE peaks are fit to determine their mean charge value which is then converted to the absolute gain.
@@
Subsequently, the energy resolution performance with various reflectors is tested. The CsI(Tl) crystal is wrapped with materials including Teflon, ESR (non-glued), and Tyvek. Other experiment setups were the same as Fig. 6b, with ${}^{22}\text{Na}$ and ${}^{137}\text{Cs}$ sources used. Measurements were conducted for 600 s with the PMT operating at 850 V. The energy resolution is directly calculated from the FWHM, and the results are listed in Table 3.
@@ III.2 Comparison and Discussion
It can be seen from Table 3 that Teflon provides a better energy resolution compared to other materials. This indicates that the properties of reflector material have a significant impact on the energy resolution. Ref. [54] measured the reflectivity of commonly used reflectors. For the tested reflectors in Table 3, the reflectivity values are 0.99, 0.985, and 0.97, respectively. It can be inferred that with the Teflon reflector, the PMT collects more photons due to its higher reflectivity leading to improved energy resolution. The energy resolution with glued ESR in the earlier experiment was worse compared to the resolution with non-glued ESR obtained later. This is due to the absorption or refraction of light as it passes through the glue. Moreover, the air gap between the non-glued reflectors and the crystal enhances the collection efficiency of scintillation photons through total internal reflection. In summary, Teflon is a promising reflector material for enhancing the energy resolution of MACE ECAL.
This study primarily relies on Geant4 simulation for the design of the MACE calorimeter, requiring verification of the reliability of various physical processes in the simulation and a comparison with experimental results. To test the reliability of the simulation results discussed in Section II, we conducted a detailed simulation of the benchmark CsI(Tl) detector using Geant4 with the same physics processes. The CsI(Tl) crystal, Teflon reflector, optical grease, and CR284 PMT volume are modeled identically to the experimental setup. At the end of the simulation, the number of produced photoelectrons is recorded in a ROOT file event by event. The energy resolution is 8.2% at 511 keV in the simulation, while the experiment result is 10.3% (see Table 3).
@@
Figure 8: The photoelectron spectra of ${}^{137}\text{Cs}$ in experiment and simulation.
The difference in energy resolution between the experiment and the simulation is reasonable. Fig. 8 shows the spectra of photoelectrons reconstructed from experimental data and those recorded in simulation using ${}^{137}\text{Cs}$ source. Considering the mean photoelectron number of the 662 keV peak, the Geant4 simulation results are generally consistent with the experiment. However, there are still minor, predictable discrepancies. Several possible explanations are as follows. Firstly, the PMT gain value measurement may contain deviations caused by electronic noise or statistical errors. As a result, the number of photoelectrons reconstructed from the waveform is biased by approximately 76 PEs compared with the simulation. Secondly, the ideal simulation did not account for the natural radioactive background, resulting in a lack of response in the higher energy region. However, this contribution is expected to be limited due to the coincidence and reconstruction.
@@
The comparison between the experiment and the simulation of the benchmark detector gives the expected results with little divergence, which validates the design of the MACE ECAL and provides a foundation for further research.
@@ IV Conclusions and Perspectives
After more than two decades of development of high-intensity muon source and detector technique, the first proposed muon-CLFV experiment in China, MACE, is expected to enhance the sensitivity of M-to-$\bar{\text{M}}$ conversion by three orders of magnitude. This study focuses on the offline-software-based design of the $4\pi$ ECAL geometry for MACE. The geometry parameters were optimized using the Monte-Carlo simulation method, and the performance such as energy resolution, signal efficiency, and background level were evaluated. Simulation results indicate an energy resolution of 9% at 511 keV, and a signal efficiency of 55.5% for annihilation $\gamma$-ray events at maximum, significantly outperforming the MACS calorimeter design. This research establishes a foundation for the subsequent technical development of the ECAL and supports the conceptual design of MACE.

### v3

@@ II.3.2 Beam-related backgrounds
A parallel simulation involving $10^{11}$ $e^{+}$ events was conducted for comparison with the signal. No coincidence events were observed in this simulation. According to the technical specifications of the Experimental Muon Source (EMuS) proposed at CSNS [22], the optimal ratio of positrons to muons is expected to be 0.1% [55]. The beam positron background level is estimated based on this data. Assuming a $\mu^{+}$ beam flux of $10^{8}/$s, and a coincidence probability of $10^{-4}$ with the magnetic spectrometer, the upper limit of the background rates is calculated to be $0.007/(10^{8}~{}\mu^{+}/\text{s}\cdot 1~{}\text{yr})$ at 90% C.L.
II.3.3 Cosmic-Ray Muon backgrounds
Cosmic-ray muons are another source of background in ECAL. The EcoMug generator [56] is integrated into the MACE offline software to simulate cosmic-ray muon events on various types of surfaces (plane, cylinder, and half-sphere) with accurate angle and momentum distributions. To evaluate cosmic-ray muon background level in a potential future environment for the MACE detector, a hypothetical 18 m deep tunnel has been modeled in Geant4. A concrete volume, represented by G4Box with dimensions $50\times 50\times 18~{}\text{m}^{3}$, is created, with an air-filled cylindrical region, represented by G4Tubs, subtracted from its bottom to form a tunnel with a 1 m radius and 10 m length. ECAL is positioned at the center of the tunnel, surrounded by a copper solenoid and an iron shield. Cosmic-ray muons are emitted from a $50\times 50~{}\text{m}^{2}$ plane source at sea level. With a rate of 129 Hz, cosmic-ray muon events equivalent to 1 hour of exposure are generated.
The simulation results indicate that cosmic-ray muons still contribute to the background in the ECAL even after passing through the concrete and shielding. This includes some events where $e^{+}e^{-}$ annihilation produces 511 keV $\gamma$ rays. After applying coincidence selection, no events remain, resulting in background rates of less than 2/2/2 /yr when considering coincidences with the Magnetic Spectrometer. Typically, a veto system can be installed outside ECAL to further reject cosmic-ray muon background [6, 57].
@@ III.1 Experiment Setups and Results
To characterize the parameters of the CsI(Tl) scintillator, a benchmark detector was constructed and calibrated using various radioactive sources at Sun Yat-sen University (SYSU). It consists of a $3\times 3\times 8~{}\text{cm}^{3}$ CsI(Tl) crystal and a 1.5-inch PMT (HAMAMATSU CR284 with E2183-500 socket assembly). The crystal is wrapped with ESR (glued) and coupled with PMT using optical grease, finally coated with 3M tape (Fig. 6b). The readout system uses a Lecroy HDO4054A oscilloscope (12 bit, 10 GS/s, see Fig. 6c) for signal triggering, integration, and digitization. The benchmark detector is calibrated using ${}^{137}\text{Cs}$ ($7.81\times 10^{3}$ Bq) and ${}^{152}\text{Eu}$ ($3.55\times 10^{4}$ Bq) radioactive sources. Radioactive sources were positioned close to the detector, and data acquisition was performed for 1800 s. When the PMT pulse signal exceeds the threshold, the oscilloscope triggers, integrates the waveform over a 2 $\mu$s time window and records the charge value. The calibration used the 662 keV peak from ${}^{137}\text{Cs}$ and the 40 keV, 122 keV, 244 keV, 344 keV, and 1408 keV peaks from ${}^{152}\text{Eu}$, as shown in Fig. 7. The spectra display the full-energy peaks of both isotopes. The energy resolution at 511 keV is calculated to be 11.5% by fitting the FWHM curve.
@@
(c) Lecroy HDO4054A displaying a typical waveform (pulse width of $\sim 2~{}\mu$s) of CsI(Tl).
@@
Figure 7: Measured spectra of ${}^{137}\text{Cs}$ (green line), ${}^{152}\text{Eu}$ (red line) at SYSU and ${}^{22}\text{Na}$ (blue line) at IMP.
A more precise experiment was conducted to investigate the energy resolution at the Institute of Modern Physics (IMP), Chinese Academy of Sciences in Lanzhou, China. The absolute gain of the CR284 PMT was accurately calibrated in advance using single photoelectron (SPE) response measurements. A waveform generator sends pulses with a voltage of 2.65 V, a frequency of 10 kHz, and a width of 5 ns to an LED to produce weak light emission. A 120 s measurement scans the high voltage ranging from 850 V to 1250 V. The PMT response waveforms are read out and recorded in ROOT files using the CAEN V1730SB digitizer (14 bit, 500 Ms/s, see Fig. 6d), which is externally triggered by the waveform generator. After offline pulse processing, the SPE peaks are fit to determine their mean charge value which is then converted to the absolute gain.
@@
Subsequently, the energy resolution performance with various reflectors is tested. The CsI(Tl) crystal is wrapped with materials including Teflon, ESR (non-glued), and Tyvek. Other experimental setups remained identical to Fig. 6b, except for the use of ${}^{22}\text{Na}$ ($1.48\times 10^{5}$ Bq) and ${}^{137}\text{Cs}$ ($4.07\times 10^{4}$ Bq) sources. Measurements were conducted for 600 s with an integration time window of $2~{}\mu$s, under 850 V operating voltage of PMT. The energy resolution is directly calculated from the FWHM, and the results are listed in Table 4.
@@ III.2 Comparison and Discussion
It can be seen from Table 4 that Teflon provides a better energy resolution compared to other materials. This indicates that the properties of reflector material have a significant impact on the energy resolution. Ref. [58] measured the reflectivity of commonly used reflectors. For the tested reflectors in Table 4, the reflectivity values are 0.99, 0.985, and 0.97, respectively. It can be inferred that with the Teflon reflector, the PMT collects more photons due to its higher reflectivity leading to improved energy resolution. The energy resolution with glued ESR in the earlier experiment was worse compared to the resolution with non-glued ESR obtained later. This is due to the absorption or refraction of light as it passes through the glue. Moreover, the air gap between the non-glued reflectors and the crystal enhances the collection efficiency of scintillation photons through total internal reflection.
(a) ${}^{22}\text{Na}$ source.
(b) ${}^{137}\text{Cs}$ source.
@@
Figure 8: The energy spectra of ${}^{22}\text{Na}$ and ${}^{137}\text{Cs}$ in experiment (dots with error bars) and simulation (red line).
This study primarily relies on Geant4 simulation for the design of the MACE calorimeter, requiring verification of the reliability of various physical processes in the simulation and a comparison with experimental results. To test the reliability of the simulation results discussed in Section II, we conducted a detailed simulation of the benchmark CsI(Tl) detector using Geant4 with the same physics processes. The CsI(Tl) crystal, Teflon reflector, optical grease, and CR284 PMT volume are modeled identically to the experimental setup. Fig. 8 shows the spectra of calibrated experimental data and those recorded in simulation using ${}^{22}\text{Na}$ and ${}^{137}\text{Cs}$ source. The energy resolution is 9.7% at 511 keV in the simulation, compared to 10.3% in the experiment (see Table 4), resulting in a relative difference of 6%. Note that the difference in the width and height of the full-energy peak between the MC and the data originates from the discrepancy in energy resolution. Other minor discrepancies in the spectra shape are mainly likely caused by the insufficient modeling of electronic noise and nonlinearity, natural backgrounds, materials around the detector, and the simplified optical processes in Geant4 [59]. This work focuses on a bench test of the CsI(Tl) crystal. However, important factors mentioned above that are difficult to simulate will be further studied in future prototype development efforts.
In addition, we tested the SiPM readout using the NDL EQR15 11-6060D-S ($6\times 6~{}\text{mm}^{2}$) with a bias voltage set to 42 V in the Teflon reflector case. Using the same approach, an energy resolution of 26.1% is obtained. This significantly worse performance is attributed to the much smaller sensitive area of a single SiPM compared to the PMT. Simulation with the same configuration indicates that using a $4\times 4$ SiPM array with a total area of $24\times 24~{}\text{mm}^{2}$ could approximately double the number of collected photoelectrons, resulting in an energy resolution of 6.58% at 511 keV.
@@
The comparison between the experiment and the simulation of the benchmark detector gives the expected results of energy resolution at 511 keV, which validates the design of the MACE ECAL and provides a foundation for further research. The SiPM array readout, with its high PDE, low voltage requirements, and tolerance to magnetic fields, represents a promising upgrade option for the future.
@@ IV Conclusions and Perspectives
After more than two decades of development of high-intensity muon source and detector technique, the first proposed muon-CLFV experiment in China, MACE, is expected to enhance the sensitivity of M-to-$\bar{\text{M}}$ conversion by more than two orders of magnitude. This study focuses on the offline-software-based design of the $4\pi$ ECAL geometry for MACE. The geometry parameters were optimized using the Monte-Carlo simulation method, and the performance such as energy resolution, signal efficiency, and background level were evaluated. Simulation results indicate an energy resolution of 10.8% at 511 keV, and a signal efficiency of 78.3% for annihilation $\gamma$-ray events, greatly outperforming the MACS calorimeter design. This research establishes a foundation for the subsequent technical development of the ECAL and supports the conceptual design of MACE.